1. The following are the values of waiting time of 6 customers of a city bank in minutes β, 5, 6, 6, 8, 9 Find mean and stanidará deviations using the formula: n(n-1)
Answer #1

Ans:

x x^2
3 9
5 25
6 36
6 36
8 64
9 81
Total 37 251

Mean=37/6=6.167

s^2=(6*251-37^2)/(6*(6-1))

s^2=4.567

standard deviation,s=sqrt(4.567)=2.137